OpenStudy (animelover98):

Use the table of the function f(x) = x4 - 2x3 to answer this question. x f(x) -1 3 0 0 1 -1 2 0 3 27 What is the average rate of change from x = -1 to x = 1? -2 0 2 4

OpenStudy (aum):

Average rate of change from x = a to x = b is given by the formula: \[ \frac{f(b)-f(a)}{b-a} \]Average rate of change from x = -1 to x = 1 is given by:\[ \frac{f(1)-f(-1)}{1-(-1)} = ? \]

OpenStudy (animelover98):

What does the f in this equation stand for? Do I choose any function of x from the chart?

OpenStudy (aum):

f(x) is the function. f(b) is the value of the function when x = b You are given a table of x values and the corresponding f(x) values: x f(x) -1 3 when x = -1, f(-1) = 3 0 0 1 -1 when x = 1, f(1) = -1 2 0 3 27 \[ \frac{f(1)-f(-1)}{1-(-1)} = \frac{-1-3}{1+1} = \frac{-4}{2} = -2 \]